# Current validation practice undermines surgical AI development

### v1

@@ Abstract.
Abstract: Surgical data science (SDS) is rapidly advancing, yet clinical adoption of artificial intelligence (AI) in surgery remains severely limited, with inadequate validation emerging as a key obstacle. In fact, existing validation practices often neglect the temporal and hierarchical structure of intraoperative videos, producing misleading, unstable, or clinically irrelevant results. In a pioneering, consensus-driven effort, we introduce the first comprehensive catalog of validation pitfalls in AI-based surgical video analysis that was derived from a multi-stage Delphi process with 91 international experts. The collected pitfalls span three categories: (1) data (e.g., incomplete annotation, spurious correlations), (2) metric selection and configuration (e.g., neglect of temporal stability, mismatch with clinical needs), and (3) aggregation and reporting (e.g., clinically uninformative aggregation, failure to account for frame dependencies in hierarchical data structures). A systematic review of surgical AI papers reveals that these pitfalls are widespread in current practice, with the majority of studies failing to account for temporal dynamics or hierarchical data structure, or relying on clinically uninformative metrics. Experiments on real surgical video datasets provide the first empirical evidence that ignoring temporal and hierarchical data structures can lead to drastic understatement of uncertainty, obscure critical failure modes, and even alter algorithm rankings. This work establishes a framework for the rigorous validation of surgical video analysis algorithms, providing a foundation for safe clinical translation, benchmarking, regulatory review, and future reporting standards in the field.
@@ 1. Main
The research field of surgical data science (SDS) was formally introduced in 2017, establishing it as a distinct field at the intersection of surgery, data science, and artificial intelligence (AI)111Throughout this manuscript, we use the term ”surgical” in a procedural sense, consistent with the definition of SDS as encompassing ”all clinical disciplines in which patient care requires intervention to manipulate anatomical structures with a diagnostic, prognostic, or therapeutic goal, such as surgery, interventional radiology, radiotherapy and interventional gastroenterology” [maier2017surgical]. This includes closely related interventional domains such as endoscopy, which share video-centric data, workflows, and validation challenges. [maier2017surgical]. The highly interdisciplinary domain leverages data acquisition, analysis, and modeling to enhance surgical decision-making, execution, training, and patient outcomes throughout the entire surgical care pathway. Since its beginnings, SDS has shown remarkable growth in AI-based publications [li2025artificial] (e.g., [hashimoto2019computer, madani2022artificial, kiyasseh2023vision, khan2024artificial]). However, clinical translation of SDS methods remains limited. For example, while over 1,000 AI-enabled medical devices have received authorization from the US Food and Drug Administration (FDA) since 2017, only six AI-enabled devices have been specifically approved by the FDA General and Plastic Surgery Devices panel, and only 17 for the field of gastroenterology-urology222https://www.fda.gov/medical-devices/software-medical-device-samd/artificial-intelligence-enabled-medical-devices as of 2025.
@@
While multiple factors contribute to the limited clinical translation of SDS methods (e.g., workflow integration or regulatory hurdles), a key bottleneck lies in the lack of robust and rigorous validation of SDS algorithms as a prerequisite for safety and efficacy. Previous works have shown that validation in AI-driven medical image analysis is often flawed [carstens2025artificial, reinke2024understanding, funke2023metrics, kofler2023we, vaassen2020evaluation]. For example, it has been shown that the choice of metrics that do not reflect the underlying biomedical research question can severely undermine the validity of validation outcomes [reinke2021common, reinke2024understanding]. To address this problem, the Metrics Reloaded initiative introduced a comprehensive list of metric-related pitfalls as well as a problem-aware metric recommendation framework guiding researchers in finding the most suitable metrics for problems related to classification (image-level classification, object detection, semantic/instance segmentation) [maier2024metrics].
@@
commonly used metrics may fail to capture temporal stability, rewarding flickering predictions with artificially high scores, and
@@
The present work aims to initiate a paradigm shift in the validation of surgical AI algorithms. It evolved from an international, multidisciplinary effort initiated at a dedicated workshop during the 2023 annual meeting of the Society of American Gastrointestinal and Endoscopic Surgeons (SAGES) – one of the leading international societies for minimally invasive surgery. The workshop brought together leading experts in SDS, surgery, and AI, and laid the foundation for translating findings from the Metrics Reloaded initiative into the video-centric context of surgery. Building on the outcomes of this workshop, we subsequently launched a large-scale Delphi process that brought together a global panel of more than 90 experts to systematically identify, refine, and consolidate validation pitfalls specific to surgical video analysis.
Specifically, this work makes the following pioneering contributions to the safe clinical adoption of surgical AI:
Consensus-based catalog of validation pitfalls: We introduce the first comprehensive list of pitfalls in the validation of surgical AI, together with their potential consequences and real-world risks. This catalog resulted from a combined approach including a literature review, agentic internet research, and a consensus-driven expert process.
Evidence for high occurrence of pitfalls: Through a systematic literature review, we demonstrate that these pitfalls frequently occur in current surgical AI studies.
Experimental demonstration of impact of pitfalls: Using real surgical data, we experimentally quantify how these pitfalls distort performance assessment and mask critical failure modes.
@@ 2. Results
Over a period of about three years, we conducted a structured process involving a total of 91 experts from surgery, computer vision, and data science across 68 institutions, ensuring a broad range of perspectives across both surgical practice and technical disciplines. A core of the initiative was a hypothesis-generating workshop with dedicated focus group discussions at the annual SAGES 2023 meeting and subsequently evolved into a multi-stage Delphi process. This iterative process enabled the development of a comprehensive theoretical foundation and a consensus-based catalog of validation pitfalls related to surgical video analysis, supported by a systematic review demonstrating their prevalence and by experiments quantifying their impact on performance assessment.
@@ 2.1. A multi-stage, multi-stakeholder Delphi process revealed numerous pitfalls in surgical AI validation
The resulting pitfall catalog is summarized in Fig. 2 and was structured into three key pitfall categories: [P1] pitfalls related to data, [P2] pitfalls related to metric selection and configuration, and [P3] pitfalls related to metric aggregation and reporting. Each category represents a distinct level at which pitfalls can compromise the validity of validation outcomes. Each pitfall was mapped to specific potential consequences, such as introduction of biases, unreliable performance assessment, or undetected failure modes, and to associated real-world risks (e.g., regulatory delay, compromised surgical safety), as illustrated in Fig. 2. Concrete definitions are provided in Tab. SN 2.1 (Supplementary Notes), with visual examples in Figs. 1, 4-6, and Extended Data Figs. 1- 9. Below, we outline the collected pitfalls identified within each category, highlighting why they are particularly critical in surgical AI validation. While some issues are known from other machine learning domains, the surgical setting adds unique layers of complexity and severity.
@@
P1.1: Non-representativeness and low relevance of data: The lack of representative data is a familiar challenge in machine learning, where diversity is key for reliable validation. In surgery, however, this problem is amplified by diverse sources of variability, including procedures varying across hospitals, operating room (OR) setups, surgical technique, variations in data quality (Extended Data Fig. 1a), pronounced geographical imbalance of available datasets (Extended Data Fig. 1b), and even surgical team dynamics.
P1.2: Limited sample size/test cohorts, or data imbalance: Sample size limitations are widely discussed in medical imaging AI (e.g., [christodoulou2025false]), but surgical data is particularly hard to collect. Many surgical procedures are not routinely recorded, and rare but safety-critical events may appear only occasionally. Even single videos can last hours and require extensive annotation [carstens2025artificial]. Moreover, the resulting datasets often exhibit substantial class imbalance and small, heterogeneous test cohorts, which can lead to unstable and unreliable performance estimates (Extended Data Fig. 2).
@@
P1.5: Unreliable or inconsistent annotation: High inter-rater variability is a common problem in medical imaging AI [kolbinger2025appendix300, joskowicz2019inter, funke2023metrics], yet surgical tasks are especially ambiguous (e.g., phase transitions, fine tool-tissue interactions). Even trained raters frequently disagree and maintaining consistency over frames of a video of multiple hours is particularly challenging (Fig. 1a).
P1.6: Poor data splitting: Data leakage is a common problem in machine learning. However, surgical videos are highly redundant, i.e., a single patient (= case) may generate thousands of dependent frames. Without strict separation between data subsets, results may measure memorization rather than true generalization to unseen procedures (Extended Data Fig. 4).
@@
P2.1: Mismatch of metrics and clinical needs: Choosing metrics that reflect the actual clinical objectives is important for every research field [maier2024metrics]. However, the gap between existing metrics and needs is particularly wide for surgical applications. Systems must operate in real time, keep latency within safe limits, and produce outputs that remain temporally stable across rapidly changing scenes. Existing measures rarely capture those aspects (Extended Data Fig. 5), and, for many clinically relevant aspects, no established metric may yet exist.
@@
P2.3: Inappropriate metric selection for handling annotation uncertainties: Ambiguous labels occur across domains, but phase boundaries or subtle tool-tissue contacts, among others, make ambiguity even more complex in surgical videos. In surgical reality, human spatiotemporal understanding is often associated with considerable inter-rater variability and inherent uncertainty [kolbinger2025appendix300]. Metrics assuming confident labels can either exaggerate or underestimate errors (Extended Data Fig. 6).
@@
P3.3: Lack of contextualization of performance values: Point estimates without context are problematic in any field, but even more so in surgical AI as they can be highly misleading (Extended Data Fig. 9a). For example, a high Accuracy may mainly reflect routine phases, while errors cluster in moments of adverse events.
@@ 2.2. Validation flaws are widespread in common practice
While pitfalls can theoretically occur in any validation study, their actual prevalence in state-of-the-art surgical AI publications remained unclear. To address this, we conducted a systematic screening of all papers at the 2023 Medical Image Computing and Computer Assisted Intervention (MICCAI) conference that applied deep learning methods to surgical data. As the leading international conference for medical image computing and computer-assisted interventions, MICCAI provides a representative overview of current practices in surgical AI. Key results of this analysis are summarized in Fig. 3 and in Suppl. Note 3.
@@
Surgical data cohorts were typically small and fragmented. The median dataset contained 37 training, 10 validation, and 22 test videos (minimum: 6 training, 2 validation, 2 test videos). 79% of datasets were only used once, with 38 distinct datasets identified across papers. Moreover, only 47% explicitly reported an untouched test set, while this was unclear in 37% of papers.
@@ 2.3. Experiments demonstrate consequences of pitfalls using real-world data
Surgical data are inherently hierarchical. Frames from a single video are not independent, as they come from the same patient and are influenced by factors such as the performing surgeon, hospital, or used surgical tools. Ignoring this structure can lead to unreliable performance estimates. As shown in the previous section, only a fraction of studies properly address hierarchical data, yet little empirical evidence exists on how this practice may affect uncertainty estimates.
To determine this impact, we analyzed two of the most widely used real-world datasets: (1) the Robust Medical Instrument Segmentation (RobustMIS) 2019 challenge dataset [ross2021comparative] for binary instrument segmentation and (2) the CholecTriplet dataset [nwoye2023cholectriplet2021] for surgical action triplet recognition. For both tasks, we compared CIs derived from a naïve bootstrap approach, which does not account for hierarchical dependencies, against CIs from a hierarchical bootstrap (Fig. 4).
Accounting for hierarchical data structure led to substantially wider CIs, which reflects the additional variance introduced at the video (i.e., patient) level that is ignored when assuming independence across all samples (naïve approach). Concretely, for binary segmentation, the CI widths increased by a median factor of more than 2 for both Dice similarity coefficient (DSC; 2.4x wider) and Normalized surface Dice (NSD; 2.2x wider). For surgical action recognition, CIs were 13.5x wider for mean Average Precision (mAP), 11x wider for weighted mAP, and 7.1x wider for top-5 Accuracy. These findings demonstrate that ignoring data dependencies can drastically understate model uncertainty, potentially giving a false sense of algorithm reliability in surgical settings.
@@
To shed light on this problem, we compared global results with stratified analysis on multi-instance instrument segmentation results from the RobustMIS challenge [ross2021comparative], using metadata describing various relevant, potentially confounding image properties [ross2023beyond] (Fig. 5). While aggregated DSC scores suggested stable performance across algorithms, stratification by clinically relevant conditions revealed considerable performance drops. For instance, the median DSC decreased by 0.36 (up to 0.51 for one algorithm) in frames with intersecting instruments, and smaller but clear declines appeared for smoke and motion artefacts. Purely reporting globally aggregated values can therefore be highly misleading, whereas stratification exposes failure cases in safety-critical situations.
@@
We analyzed results from the binary segmentation task of the RobustMIS challenge [ross2021comparative] and applied six different aggregation strategies: frame-wise, video-wise, phase-wise, phase-wise video-wise, video-wise phase-wise, and weighted phase-wise (see Methods (Sec. 4) for detailed descriptions of each strategy). We then compared the results for each strategy with the default frame-wise aggregation (Fig. 6). Similarly to the original challenge, we used the 5% percentile as the aggregation operator to reflect worst-case performance.
The median Kendall’s tau correlation [kendall1938new] across the different rankings compared to the default ranking was 0.68, indicating high variance in the leaderboards. The original winner changed in 80% of rankings, with a median absolute rank change of 1 and a maximum change of 3. Negative rank shifts occurred in 58% of cases, positive in 28%, and no change in 14%. As shown in Fig. 6b, when performance differences between algorithms were modest, even small changes in the aggregation led to substantial ranking shifts, questioning the reliability of the winner. These findings show that minor reporting omissions such as unspecified aggregation can substantially affect ranking conclusions and potentially influence which algorithms are prioritized for clinical translation.

### v2

@@ Abstract.
Abstract: Surgical data science (SDS) is rapidly advancing, yet clinical adoption of artificial intelligence (AI) in surgery remains limited, with inadequate validation emerging as an important contributing factor. In fact, existing validation practices often neglect the temporal and hierarchical structure of intraoperative videos, producing misleading, unstable, or clinically irrelevant results. In a pioneering, consensus-driven effort, we introduce a comprehensive catalog of validation pitfalls in AI-based surgical video analysis that was derived from a multi-stage Delphi process with 92 international experts. The collected pitfalls span three categories: (1) data (e.g., incomplete annotation, spurious correlations), (2) metric selection and configuration (e.g., neglect of temporal stability, mismatch with clinical needs), and (3) aggregation and reporting (e.g., clinically uninformative aggregation, failure to account for frame dependencies in hierarchical data structures). A systematic review of surgical AI papers reveals that these pitfalls are widespread in current practice, with the majority of studies failing to account for temporal dynamics or hierarchical data structure, or relying on clinically uninformative metrics. Experiments on real surgical video datasets provide empirical evidence that ignoring temporal and hierarchical data structures can substantially understate uncertainty, obscure critical failure modes, and even alter algorithm rankings. To address these shortcomings, we provide a catalogue of best practices compiled in a multi-stage Delphi process. Together, this work provides an evidence-based framework to inform more rigorous validation of surgical video analysis algorithms and to guide future efforts in benchmarking, reporting, regulatory review, and clinical translation.
@@ 1. Main
The research field of surgical data science (SDS) was formally introduced in 2017, establishing it as a distinct field at the intersection of surgery, data science, and artificial intelligence (AI)111Throughout this manuscript, we use the term ”surgical” in a procedural sense, consistent with the definition of SDS as encompassing ”all clinical disciplines in which patient care requires intervention to manipulate anatomical structures with a diagnostic, prognostic, or therapeutic goal, such as surgery, interventional radiology, radiotherapy, and interventional gastroenterology” [Maier-Hein et al., 2017]. This includes closely related interventional domains such as endoscopy, which share video-centric data, workflows, and validation challenges. [Maier-Hein et al., 2017]. The highly interdisciplinary domain leverages data acquisition, analysis, and modeling to enhance surgical decision-making, execution, training, and patient outcomes throughout the entire surgical care pathway. Since its beginnings, SDS has shown remarkable growth in AI-based publications [Li et al., 2025] (e.g., [Hashimoto et al., 2019, Madani et al., 2022, Kiyasseh et al., 2023, Khan et al., 2024]). However, clinical translation of SDS methods remains limited. For example, while over 1,000 AI-enabled medical devices have received authorization from the US Food and Drug Administration (FDA) since 2017, only six AI-enabled devices have been specifically approved by the FDA General and Plastic Surgery Devices panel, and only 17 for the field of gastroenterology-urology222https://www.fda.gov/medical-devices/software-medical-device-samd/artificial-intelligence-enabled-medical-devices as of 2025.
Multiple barriers persist that currently limit the clinical translation of SDS methods. These mainly include regulatory and economic hurdles as well as issues pertaining to workflow and hardware integration and clinical utility. Although barriers are manifold, a key cross-cutting concern can be observed in the lack of robust and rigorous validation of AI algorithms [Bürger et al., 2024, Madai and Higgins, 2021, Chouffani El Fassi et al., 2024, Gong et al., 2025, Bernstam et al., 2022, Olaye and Seixas, 2023, Carstens et al., 2025]. While large-scale failures of surgical AI are rarely reported, likely reflecting their still limited clinical deployment, several high-profile cases from more mature domains such as radiology and clinical decision support illustrate that shortcomings in validation have interacted with, and in some cases exacerbated, these barriers, contributing to failures during real-world deployment. For instance, the pioneering clinical decision support system IBM Watson for Oncology was validated primarily on narrow, curated cases from a single institution, which limited its ability to generalize to heterogeneous real-world patient populations, ultimately undermining clinical safety and economic viability [Strickland, 2019]. Google Health’s diabetic retinopathy system achieved high retrospective accuracy, yet deployment revealed unvalidated workflow constraints (e.g., image acquisition conditions and throughput), resulting in reduced usability and staff frustration rather than clinical benefit [Beede et al., 2020]. In another case, the Epic Sepsis Model, deployed based on proprietary internal validation, showed substantial performance degradation under independent external validation, raising concerns about patient safety and regulatory oversight [Wong et al., 2021]. An example from surgical AI highlights related limitations. Computer-aided detection systems for polyp detection were primarily validated on improved detection metrics, yet a randomized trial in community-based practice showed no improvement in clinically relevant adenoma detection, while increasing the detection of non-relevant findings and prolonging procedure time, thereby limiting clinical utility [Wei et al., 2022, Berzin and Brown, 2023].
These examples illustrate that validation limited to retrospective accuracy or narrow datasets fails to capture clinical utility, robustness, generalizability, and workflow impact, which are key determinants of real-world adoption. In line with this perspective, policy reports and reviews identify insufficient validation protocols as one of several recurring challenges affecting the deployment of AI in healthcare [for Health et al., 2025]. A high profile opinion piece in Nature Medicine went as far as to call out a “benchmarking crisis” across the entire field of biomedical machine learning, highlighting the lack of standardized validation as an important bottleneck impeding progress and patient benefit [Mahmood, 2025].
@@
Across domains, validation outcomes strongly depend on how clinical objectives are translated into quantitative assessment, in particular through the choice of performance metrics. Previous work has shown that inappropriate metrics can compromise the validity in AI-driven medical image analysis is often flawed [Carstens et al., 2025, Reinke et al., 2024, Funke et al., 2023, Kofler et al., 2023, Vaassen et al., 2020]. For example, it has been shown that the choice of metrics that do not reflect the underlying biomedical research question can severely undermine the validity of validation outcomes [Reinke et al., 2021, 2024]. To address this problem, the global Metrics Reloaded initiative introduced a comprehensive list of metric-related pitfalls as well as a problem-aware metric recommendation framework guiding researchers in finding the most suitable metrics for problems related to classification (image-level classification, object detection, semantic/instance segmentation) [Maier-Hein et al., 2024].
@@
commonly used metrics may fail to capture temporal stability and therefore not distinguish between temporally stable and rapidly fluctuating (‘flickering’) predictions, which can be perceived as irritating or implausible in clinical workflows, and
@@
The present work aims to inform a shift toward more rigorous and clinically grounded validation of surgical AI algorithms. Its initiating event was a dedicated workshop during the 2023 annual meeting of the Society of American Gastrointestinal and Endoscopic Surgeons (SAGES) – one of the leading international societies for minimally invasive surgery. The workshop brought together leading experts in SDS, surgery, and AI, and laid the foundation for translating findings from the Metrics Reloaded initiative into the video-centric context of surgery. Building on the outcomes of this workshop, over the past three years, we conducted a large-scale Delphi process that brought together a global panel of more than 90 experts to systematically identify, refine, and consolidate validation pitfalls specific to surgical video analysis.
Specifically, this work makes the following contributions to the safe clinical adoption of surgical AI:
Consensus-based catalog of validation pitfalls: We introduce a comprehensive list and structured single point of access for pitfalls in the validation of surgical AI. The catalogue is organized along the validation pipeline (data, metric selection and configuration, metric aggregation and reporting) and links each pitfall to their potential consequences and real-world risks. It resulted from a combined approach including a literature review, agentic internet research, and a consensus-driven expert process.
Evidence for high occurrence of pitfalls: Through a systematic literature review, we provide empirical evidence that these pitfalls frequently occur in current surgical AI studies, including widespread reliance on frame-wise validation and naive aggregation strategies that ignore the temporal and hierarchical structure of surgical video data.
Experimental demonstration of impact of pitfalls: Using real surgical data, we experimentally quantify how these pitfalls distort performance assessment, mask critical failure modes, and can substantially affect uncertainty estimates and algorithm rankings.
Practical, pitfall-specific best practices: Building on the identified pitfalls and supporting evidence, we derive expert-consensus best practices to support more robust validation practice and reporting in surgical AI based on a multi-stage Delphi process.
@@ 2. Results
Over a period of about three years, we conducted a structured process involving a total of 92 experts from surgery, computer vision, and data science across 68 institutions, ensuring a broad range of perspectives across both surgical practice and technical disciplines. A core of the initiative was a hypothesis-generating workshop with dedicated focus group discussions at the annual SAGES 2023 meeting and subsequently evolved into a multi-stage Delphi process. This iterative process enabled the development of a comprehensive theoretical foundation and a consensus-based catalog of validation pitfalls related to surgical video analysis, supported by a systematic review demonstrating their prevalence and by experiments quantifying their impact on performance assessment.
@@ 2.1. A multi-stage, multi-stakeholder Delphi process revealed numerous pitfalls in surgical AI validation
The resulting pitfall catalog is summarized in Fig. 2 and was structured into three key pitfall categories: [P1] pitfalls related to data, [P2] pitfalls related to metric selection and configuration, and [P3] pitfalls related to metric aggregation and reporting. Each category represents a distinct level at which pitfalls can compromise the validity of validation outcomes. Each pitfall was mapped to specific potential consequences, such as introduction of biases, unreliable performance assessment, or undetected failure modes, and to associated real-world risks (e.g., regulatory delay, compromised surgical safety), as illustrated in Fig. 2. Concrete definitions are provided in Tab. SN 1.1 (Supplementary Notes), with visual examples in Figs. 1, 4-6, and Extended Data Figs. 1-9. Below, we outline the collected pitfalls identified within each category, highlighting why they are particularly critical in surgical AI validation. While some issues are known from other machine learning domains, the surgical setting adds unique layers of complexity and severity.
@@
P1.1: Non-representativeness and low relevance of data: The lack of representative data is a familiar challenge in machine learning, where diversity is key for reliable validation. In surgery, however, this problem is amplified by diverse sources of variability, including procedures varying across hospitals, operating room (OR) setups, surgical technique, variations in data quality (Extended Data Fig. 1a), pronounced geographical imbalance of available datasets (Extended Data Fig. 1b), and heterogeneity in surgical team composition and experience levels.
P1.2: Limited sample size/test cohorts, or data imbalance: Sample size limitations are widely discussed in medical imaging AI (e.g., [Christodoulou et al., 2025]), but surgical data is particularly hard to collect. Many surgical procedures are not routinely recorded, and rare, but safety-critical events may appear only occasionally. Even single videos can last hours and require extensive annotation [Carstens et al., 2025]. Moreover, the resulting datasets often exhibit substantial class imbalance and small, heterogeneous test cohorts, which can lead to unstable and unreliable performance estimates (Extended Data Fig. 2).
@@
P1.5: Unreliable or inconsistent annotation: High inter-rater variability is a common problem in medical imaging [Kolbinger et al., 2025, Joskowicz et al., 2019, Funke et al., 2023], yet surgical tasks are especially ambiguous (e.g., phase transitions, fine tool-tissue interactions). Even trained raters frequently disagree and maintaining consistency over frames of a video of multiple hours is particularly challenging (Fig. 1a).
P1.6: Poor data splitting: Data leakage is a common problem in machine learning. Surgical videos are highly redundant, i.e., a single patient (= case) may generate thousands of dependent frames. Without strict separation between data subsets, results may measure memorization rather than true generalization to unseen procedures (Extended Data Fig. 4).
@@
P2.1: Mismatch of metrics and clinical needs: Choosing metrics that reflect the actual clinical objectives is important for every research field [Maier-Hein et al., 2024]. However, the gap between existing metrics and needs is particularly wide for surgical applications. Systems must operate in real time, keep latency within safe limits, and produce outputs that remain temporally stable across rapidly changing scenes. Existing measures rarely capture those aspects (Extended Data Fig. 5), and, for many clinically relevant aspects, no established metric may yet exist.
@@
P2.3: Inappropriate metric selection for handling annotation uncertainties: Ambiguous labels occur across domains, but phase boundaries or subtle tool-tissue contacts, among others, make ambiguity even more complex in surgical videos. In surgical reality, human spatiotemporal understanding is often associated with considerable inter-rater variability and inherent uncertainty [Kolbinger et al., 2025]. Metrics assuming confident labels can either exaggerate or underestimate errors (Extended Data Fig. 6).
@@
P3.3: Lack of contextualization of performance values: Point estimates without context are problematic in any field, but even more so in surgical AI, as they can be highly misleading (Extended Data Fig. 9a). For example, a high Accuracy may mainly reflect routine phases, while errors cluster in moments of adverse events.
@@ 2.2. Validation flaws are widespread in common practice
While pitfalls can theoretically occur in any validation study, their actual prevalence in state-of-the-art surgical AI publications remained unclear. To address this, we conducted a systematic screening of all papers at the 2023 Medical Image Computing and Computer Assisted Intervention (MICCAI) conference that applied deep learning methods to surgical data. As the leading international conference for the theme, MICCAI provides a representative overview of current practices in surgical AI. Key results of this analysis are summarized in Fig. 3 and in SUPPL. NOTE 4.
@@
Surgical data cohorts were typically small and fragmented. The median dataset contained 37 training, 10 validation, and 22 test videos (minimum: 6 training, 2 validation, 2 test videos). 79% of datasets were only used once, with 38 distinct datasets identified across papers. Moreover, only 47% explicitly reported an untouched test set, while this was unclear in 37% of the papers.
@@
To assess the robustness of our findings over time, we conducted an additional targeted screening of MICCAI 2025 SDS papers (n = 89) using a simplified screening protocol (see Methods Sec. 4 for details). Overall, results were largely consistent with those reported above. Differences were generally small, with the majority of values differing by only several percentage points, and no systematic trend towards improved validation practices. A comparison of results from the 2023 and 2025 screenings is provided in SUPPL. NOTE 5.
Notable deviations were observed for a small number of parameters. The proportion of papers with unclear reporting of an untouched test set increased substantially (from 37% to 74%), and the proportion of metrics used only once also increased markedly. Other characteristics, including dataset sizes and general reporting patterns, remained broadly comparable.
@@ 2.3. Experiments demonstrate consequences of pitfalls using real-world data
Surgical data are inherently hierarchical. Frames from a single video are not independent, as they come from the same patient and are influenced by factors such as the performing surgeon, the hospital, or the used surgical tools. Ignoring this structure can lead to unreliable performance estimates and highly misleading uncertainty quantification. Confidence intervals (CIs) are a key mechanism for characterizing model uncertainty and are explicitly required for assessing model reliability and clinical readiness in medical imaging AI, including under U.S. Food and Drug Administration (FDA) regulatory frameworks [US Food and Drug Administration et al., 2019, 2007, CFR, 2020a, b, Christodoulou et al., 2024]. As shown in the previous section, only a fraction of studies properly addresses hierarchical data, yet little empirical evidence exists on how this practice may affect uncertainty estimates such as CIs under realistic data dependencies.
To determine this impact, we analyzed two widely used real-world datasets: (1) the Robust Medical Instrument Segmentation (RobustMIS) 2019 challenge dataset [Roß et al., 2021] for binary instrument segmentation, which is a large-scale benchmark dataset with multi-team predictions, and (2) the CholecTriplet dataset [Nwoye et al., 2023a] for surgical action triplet recognition, which is one of the most frequently used SDS datasets [Carstens et al., 2025]. For both tasks, we compared CIs derived from a naïve bootstrap approach, which does not account for hierarchical dependencies, against CIs from a hierarchical bootstrap (Fig. 4).
Accounting for hierarchical data structure led to substantially wider CIs, which reflects the additional variance introduced at the video (i.e., patient) level that is ignored when assuming independence across all samples (naïve approach). Concretely, for binary segmentation, the CI widths increased by a median factor of more than 2 for both Dice similarity coefficient (DSC; 2.3x wider) and Normalized surface Dice (NSD; 2.1x wider). For surgical action recognition, CIs were 13.3x wider for mean Average Precision (mAP), 10.9x wider for weighted mAP, and 7.2x wider for top-5 Accuracy. These findings demonstrate that ignoring data dependencies can substantially understate model uncertainty, potentially giving a false sense of algorithm reliability in surgical settings.
@@
To shed light on this problem, we compared global results with stratified analysis on multi-instance instrument segmentation results from the RobustMIS challenge [Roß et al., 2021], using metadata describing various relevant, potentially confounding image properties [Roß et al., 2023] (Fig. 5). While aggregated DSC scores suggested stable performance across algorithms, stratification by clinically relevant conditions revealed considerable performance drops. For instance, the median DSC decreased by 0.34 (up to 0.52 for one algorithm) in frames with intersecting instruments, and smaller but clear declines appeared for smoke and motion artefacts. Purely reporting globally aggregated values can therefore be highly misleading, whereas stratification exposes failure cases in safety-critical situations.
@@
We analyzed results from the binary segmentation task of the RobustMIS challenge [Roß et al., 2021] and applied six different aggregation strategies: frame-wise, video-wise, phase-wise, phase-wise video-wise, video-wise phase-wise, and weighted phase-wise (see Methods (Sec. 4) for detailed descriptions of each strategy). We then compared the results for each strategy with the default frame-wise aggregation (Fig. 6). Similarly to the original challenge, we used the 5% percentile as the aggregation operator to reflect worst-case performance.
The median Kendall’s tau correlation [Kendall, 1938] across the different rankings compared to the default ranking was 0.68, indicating high variance in the leaderboards. The original winner changed in 80% of rankings, with a median absolute rank change of 1 and a maximum change of 3. Negative rank shifts occurred in 58% of cases, positive in 28%, and no change in 14%. As shown in Fig. 6b, when performance differences between algorithms were modest, even small changes in the aggregation led to substantial ranking shifts, questioning the reliability of the winner. These findings show that minor reporting omissions such as unspecified aggregation can substantially affect ranking conclusions and potentially influence which algorithms are prioritized for clinical translation.
Practical best practices to mitigate validation pitfalls
Building on the identified pitfalls and their empirical consequences, the Delphi consortium derived practical, pitfall-specific best practices, reflecting expert consensus across surgical data science. These best practices are designed to support more robust validation practice in surgical video AI by distinguishing between measures to avoid the occurrence of specific pitfalls as well as strategies to appropriately deal with them when these cannot be fully avoided, for example due to data, annotation, or workflow constraints. The best practices are intended to provide a comprehensive, structured overview of methodologically sound validation approaches identified through the Delphi process, acknowledging that their implementation depends on the specific study context and available resources.
The full list of best practices includes 124 items and is provided in SUPPL. NOTE 3. Across best practices, recurring structural patterns were identified, including early specification of validation design choices, alignment with intended clinical use, explicit consideration of variability, respect for hierarchical data structure, restriction of conclusions to the validated setting, and transparent reporting of decisions (see Fig. 7).